# Online Feature Updates Improve Online (Generalized) Label Shift Adaptation

### v1

@@ Abstract
This paper addresses the prevalent issue of label shift in an online setting with missing labels, where data distributions change over time and obtaining timely labels is challenging. While existing methods primarily focus on adjusting or updating the final layer of a pre-trained classifier, we explore the untapped potential of enhancing feature representations using unlabeled data at test-time. Our novel method, Online Label Shift adaptation with Online Feature Updates (OLS-OFU), leverages self-supervised learning to refine the feature extraction process, thereby improving the prediction model. Theoretical analyses confirm that OLS-OFU reduces algorithmic regret by capitalizing on self-supervised learning for feature refinement. Empirical studies on various datasets, under both online label shift and generalized label shift conditions, underscore the effectiveness and robustness of OLS-OFU, especially in cases of domain shifts.
@@ 1 Introduction
The effectiveness of most supervised learning models relies on a key assumption that the training data and test data share the same distribution. However, this assumption rarely holds in real-world scenarios, leading to the phenomenon of distribution shift (Quiñonero-Candela et al., 2009; Amodei et al., 2016). Previous research has primarily focused on understanding distribution shifts in offline or batch settings, where a single shift occurs between the training and test distributions (Lin et al., 2002; Shimodaira, 2000; Zadrozny, 2004; Zhang et al., 2013; Lipton et al., 2018). In contrast, real-world applications often involve test data arriving in an online fashion, and the distribution shift can continuously evolve over time. Additionally, there is another challenging issue of missing and delayed feedback labels, stemming from the online setup, where gathering labels for the streaming data in a timely manner becomes a challenging task.
@@
To tackle the distribution shift problem, prior work often relies on additional assumptions regarding the nature of the shift, such as label shift or covariate shift (Schölkopf et al., 2012). In this paper, we focus on the common (generalized) label shift problem in an online setting with missing labels (Wu et al., 2021) . Specifically, the learner is given a fixed set of labeled training data $D_{0}\sim\mathcal{P}^{\rm train}$ in advance and trains a model $f_{0}$. During test-time, only a small batch of unlabelled test data $S_{t}\sim\mathcal{P}^{\rm test}_{t}$ arrives in an online fashion ($t=1,2,\cdots$). For the online label shift, we assume the label distribution $\mathcal{P}^{\rm test}_{t}(y)$ may change over time $t$ while the conditional distribution remains the same, i.e. $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$. For example, employing MRI image classifiers for concussion detection becomes challenging as label shifts emerge from seasonal variations in the image distribution. A classifier trained during skiing season may perform poorly when tested later, given the continuous change in image distribution between skiing and non-skiing seasons. In contrast to label shift, the generalized label shift relaxes the assumption of an unchanged conditional distribution on $x$ given $y$. Instead, it assumes that there exists a transformation $h$ of the covariate, such that the conditional distribution $\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y)$ stays the same. Reiterating our example, consider an MRI image classifier that undergoes training and testing at different clinics, each equipped with MRI machines of varying hardware and software versions. As a result, the images may display disparities in brightness, resolution, and other characteristics. However, a feature extractor $h$ exists, capable of mapping these variations to the same point in the transformed feature space, such that the conditional distribution $\mathcal{P}(h(x)|y)$ remains the same. In both settings, the goal of the learner is to adapt to the (generalized) label shift within the non-stationary environment, continually adjusting the model’s predictions in real-time.
@@
Existing algorithms for online label shift adaptation (OLS) primarily adopt one of two approaches: either directly reweighting of the pretrained classifier $f_{0}$, or re-training only the final linear layer of $f_{0}$ — while keeping the feature extractor frozen. Recent work (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) have demonstrated the potential for improving feature extractors, even during test-time and in the absence of labeled data. We hypothesize that a similar effect can be harnessed in the context of (generalized) label shift, leading to the idea of improving feature representation learning during testing. In online label shift, updating the feature extractor offers two potential advantages. First, it utilizes the additional unlabeled samples, hence enhancing the sample efficiency of the feature extractor. Second, it enables the feature extractor to adapt to label shift, which is crucial to the learning process as the optimal feature extractor may depend on the underlying label distribution. Particularly in generalized label shift scenarios, where the feature transformation $h$ is often unknown, the integration of extra unlabeled test samples facilitates the learning of $h$.
@@
Building upon this insight, this paper introduces the Online Label Shift adaptation with Online Feature Updates (OLS-OFU) framework, aimed at enhancing feature representation learning in the context of online label shift adaptation. Specifically, each instantiation of OLS-OFU incorporates a self-supervised learning method associated with a loss function denoted as $l_{\rm ssl}$ for feature representation learning, and an online label shift adaptation (OLS) algorithm to effectively address distribution shift. However, it is worth pointing out that naively integrating feature extractor updates into traditional OLS algorithms will violate the underlying key assumptions of OLS algorithms (as we show in Section 3.1), hence compromising both performance and theoretical guarantees. To address this, our OLS-OFU is carefully designed to facilitate online feature updates within the OLS paradigm. At each time step, OLS-OFU executes a modified OLS algorithm, updates the feature extractor through self-supervised learning, and subsequently refines the final linear layer. This approach enables seamless online feature updates while preserving the theoretical guarantees inherent to the original OLS algorithms.
@@
In addition to its ease of implementation and seamless integration with existing OLS algorithms, OLS-OFU also shows strong theoretical guarantee and empirical performance. Theoretically, we demonstrate that OLS-OFU effectively reduces the loss of the overall algorithm by leveraging self-supervised learning techniques to enhance the feature extractor, thereby improving predictions for test samples at each time step $t$. Empirical evaluations on both CIFAR-10 and CIFAR-10C datasets, considering both online label shift and online generalized label shift scenarios, validate the effectiveness of OLS-OFU. It consistently outperforms its counterpart OLS methods across various datasets and distribution shift settings, underscoring its robustness across different scenarios.
@@ 2 Problem Setting & Related Work
We start with some basic notations. Let $\Delta^{K-1}$ be the probability simplex. Let $f:\mathcal{X}\to\Delta^{K-1}$ denote a classifier. Given an input $x$ from domain $\mathcal{X}$, $f(x)$ outputs a probabilistic prediction over $K$ classes. For example, $f$ can be the output from the softmax operation after any neural network. If we reweight a model $f$ by a vector $p\in\mathbb{R}^{K}$, we refer to this model as $g(\cdot;f,p)$ with $g$ denotes the method of reweighting. For any two vectors $p$ and $q$, $p/q$ denotes the element-wise division.
Online distribution shift adaptation. The effectiveness of any machine learning model $f$ relies on a common assumption that the train data $D_{0}$ and test data $D_{\text{test}}$ are sampled from the same distribution, i.e., $\mathcal{P}^{\rm train}=\mathcal{P}^{\rm test}$. However, this assumption is often violated in practice, which leads to distribution shift. This can be caused by various factors, such as data collection bias and changes in the data generation process. Moreover, once a well-trained model $f_{0}$ is deployed in the real world, it moves into the testing phase, which is composed of a sequence of time periods or time steps. The test distribution at time step $t$, $\mathcal{P}_{t}^{\rm test}$, from which test data $x_{t}$ is sampled, may vary over time. One example is that an MRI image classifier might be trained on MRI images collected during skiing season (which may have a high frequency of head concussions) but tested afterward (when the frequency of concussion is lower). The test stage can last several months until the next classifier is trained. During this test period, the distribution of MRI images may undergo continuous changes, transitioning between non-skiing and skiing seasons.
@@
If we have knowledge of $\mathcal{P}^{\rm test}_{t}$ or have enough samples from $\mathcal{P}^{\rm test}_{t}$, then the problem reduces to offline distribution shift for each single time step $t$. Alternatively, if we have a few labeled samples, the problem can be treated as classical online learning, and we can run stochastic gradient descent at the end of every time step $t$ to meet certain theoretical guarantees. However, dealing with very few unlabeled samples or even a single sample $x_{t}$ is more realistic yet technically highly non-trivial. An effective algorithm under this scenario has to utilize information from all historical data (train data $D_{0}$, validation data $D_{0}^{\prime}$ and test data up to time $t$), as well as previously deployed models $f_{0,1,\cdots,t-1}$. In this paper, we consider the challenging scenario where at each time step $t$, only a small batch of unlabeled samples $S_{t}=\{x_{t}^{1},\cdots,x_{t}^{B}\}$ is received. We formalize the algorithm $\mathcal{A}$ as: $\forall t\in[T]$,
@@
In contrast to the classical online learning setup, this scenario presents a significant challenge as classical online learning literature usually relies on having access to either full or partial knowledge of the loss at each time step, i.e., $\ell_{\rm sup}(f(x_{t}),y_{t})$. In this setting, however, only a batch of unlabeled samples is provided at each time step. This lack of access to label information and loss values presents a significant challenge in accurately estimating the true loss defined in Equation 1.
Related work have studied the online distribution shift adaptation under various distribution shift assumptions, such as label shift (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) and covariate shift (Zhang et al., 2023). In this paper, we focus on online (generlized) label shift adaptation.
Online label shift adaptation. Online label shift assumes
$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y),$
@@
while the marginal distribution of the label $\mathcal{P}^{\rm test}_{t}(y)$ changes over time. This assumption is most typical when the label $y$ is the causal variable and and the feature $x$ is the observation  (Schölkopf et al., 2012). The aforementioned example of concussion detection from MRI images fits this scenario, where the presence or absence of a concussion (label) causes the observed MRI image features. Most previous methods tackle this problem through a non-trivial reduction to the classical online learning problem. Consequently, most of the online label shift algorithms (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) study the theoretical guarantee of the algorithm via the convergence of the regret function, which could be either static regret or dynamic regret. In previous studies, the hypothesis class $\mathcal{F}$ of the prediction function $f$ is typically chosen in one of two ways. The first approach involves defining $\mathcal{F}$ as a family of post-hoc reweightings of $f_{0}$, with the parameter space comprising potential reweight vectors. Examples within this category include ROGD (Wu et al., 2021), FTH (Wu et al., 2021), and FLHFTL (Baby et al., 2023). The second approach defines $\mathcal{F}$ as a family of functions that share the same parameters in $f_{0}$ except the last linear layer, such as UOGD (Bai et al., 2022) and ATLAS (Bai et al., 2022).
@@
Online generalized label shift adaptation. In the context of MRI image classification, where head MRI images serve as the feature $x$, variations in software or hardware across different clinics’ MRI machines can introduce discrepancies in image characteristics like brightness, contrast, and resolution. In such scenarios, the conditional probability distribution $\mathcal{P}(x|y)$ is no longer invariant. However, when a feature extractor $h$ is robust enough, it can map images into a feature space where the images from different machines have the same distributions $\mathcal{P}(h(x)|y)$ in the transformed feature space. The concept of generalized label shift, as introduced in Tachet des Combes et al. (2020), formalizes this by postulating the existence of an unknown function $h$, such that $\mathcal{P}(h(x)|y)$ remains invariant. The primary challenge in this context is to find the underlying transformation $h$. Building upon this, online generalized label shift assumes that, there exists an unknown function $h$ such that
@@
Self-supervised learning. Inspired by the extensive body of work in semi-supervised learning (Grandvalet & Bengio, 2004b; Lee et al., 2013; Laine & Aila, 2016; Gidaris et al., 2018; Miyato et al., 2018) and unsupervised representation learning (Chen et al., 2020a; He et al., 2020; Grill et al., 2020; He et al., 2022), self-supervised learning (SSL) techniques have emerged as promising tools for improving feature extraction from unlabeled data, such as applications in image classification. Likewise, when dealing with online label shift adaptation, it is crucial to leverage the unlabeled test samples $S_{1}\cup\cdots\cup S_{t-1}$ obtained from previous time steps. Ideally, these unlabeled samples could improve the feature representation learning and ultimately lead to better prediction for the test samples $S_{t}$ at time step $t$.
In this paper, we narrow our focus on three particular SSL techniques in the evaluation for classification tasks: rotation degree prediction (Gidaris et al., 2018; Sun et al., 2020), entropy minimization (Grandvalet & Bengio, 2004b; Wang et al., 2020) and MoCo (He et al., 2020; Chen et al., 2020b, 2021). It is important to note that this concept extends beyond these three SSL techniques, and the incorporation of more advanced SSL techniques has the potential to further elevate the performance. Specifically, rotation degree prediction involves initially rotating a given image by a specific degree from the set $\{0,90,180,270\}$ and the classifier is required to determine the degree by which the image has been rotated. Entropy minimization utilizes a minimum entropy regularizer, with the motivation that unlabeled examples are mostly beneficial when classes have a small overlap. MoCo is a more advanced representation learning technique, using a query and momentum encoder to learn representations from unlabeled data by maximizing the similarity of positive pairs and minimizing the similarity of negative pairs. More details of these SSL techniques are given in Appendix E.8.

### v3

@@ Abstract
This paper addresses the prevalent issue of label shift in an online setting with missing labels, where data distributions change over time and obtaining timely labels is challenging. While existing methods primarily focus on adjusting or updating the final layer of a pre-trained classifier, we explore the untapped potential of enhancing feature representations using unlabeled data at test-time. Our novel method, Online Label Shift adaptation with Online Feature Updates (OLS-OFU), leverages self-supervised learning to refine the feature extraction process, thereby improving the prediction model. By carefully designing the algorithm, theoretically OLS-OFU maintains the similar online regret convergence to the results in the literature while taking the improved features into account. Empirically, it achieves substantial improvements over existing methods, which is as significant as the gains existing methods have over the baseline (i.e., without distribution shift adaptations).
@@ 1 Introduction
The effectiveness of most supervised learning models relies on a key assumption that the training data and test data share the same distribution. However, this assumption rarely holds in real-world scenarios, leading to the phenomenon of distribution shift [41, 2]. Previous research has primarily focused on understanding distribution shifts in offline or batch settings, where a single shift occurs between the training and test distributions [33, 45, 51, 52, 35]. In contrast, real-world applications often involve test data arriving in an online fashion, and the distribution shift can continuously evolve over time. Additionally, there is another challenging issue of missing and delayed feedback labels, stemming from the online setup, where gathering labels for the streaming data in a timely manner becomes a challenging task.
@@
To tackle the distribution shift problem, prior work often relies on additional assumptions regarding the nature of the shift, such as label shift or covariate shift [43]. In this paper, we focus on the common (generalized) label shift problem in an online setting with missing labels [49] . Specifically, the learner is given a fixed set of labeled training data $D_{0}\sim\mathcal{P}^{\rm train}$ in advance and trains a model $f_{0}$. During test-time, only a small batch of unlabelled test data $S_{t}\sim\mathcal{P}^{\rm test}_{t}$ arrives in an online fashion ($t=1,2,\cdots$). For the online label shift, we assume the label distribution $\mathcal{P}^{\rm test}_{t}(y)$ may change over time $t$ while the conditional distribution remains the same, i.e. $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$. For example, employing MRI image classifiers for concussion detection becomes challenging as label shifts emerge from seasonal variations in the image distribution. A classifier trained during skiing season may perform poorly when tested later, given the continuous change in image distribution between skiing and non-skiing seasons. In contrast to label shift, the generalized label shift relaxes the assumption of an unchanged conditional distribution on $x$ given $y$. Instead, it assumes that there exists a transformation $h$ of the covariate, such that the conditional distribution $\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y)$ stays the same. Reiterating our example, consider an MRI image classifier that undergoes training and testing at different clinics, each equipped with MRI machines of varying hardware and software versions. As a result, the images may display disparities in brightness, resolution, and other characteristics. However, a feature extractor $h$ exists, capable of mapping these variations to the same point in the transformed feature space, such that the conditional distribution $\mathcal{P}(h(x)|y)$ remains the same. In both settings, the goal of the learner is to adapt to the (generalized) label shift within the non-stationary environment, continually adjusting the model’s predictions in real-time.
@@
Existing algorithms for online label shift adaptation (OLS) primarily adopt one of two approaches: either directly reweighting of the pretrained classifier $f_{0}$, or re-training only the final linear layer of $f_{0}$ — while keeping the feature extractor frozen. Recent work [46, 48, 36, 39] have demonstrated the potential for improving feature extractors, even during test-time and in the absence of labeled data. We hypothesize that a similar effect can be harnessed in the context of (generalized) label shift, leading to the idea of improving feature representation learning during testing. In online label shift, updating the feature extractor offers two potential advantages. First, it utilizes the additional unlabeled samples, hence enhancing the sample efficiency of the feature extractor. Second, it enables the feature extractor to adapt to label shift, which is crucial to the learning process as the optimal feature extractor may depend on the underlying label distribution. Particularly in generalized label shift scenarios, where the feature transformation $h$ is often unknown, the integration of extra unlabeled test samples facilitates the learning of $h$.
@@
Building upon this insight, this paper introduces the Online Label Shift adaptation with Online Feature Updates (OLS-OFU) framework, aimed at enhancing feature representation learning in the context of online label shift adaptation. Specifically, each instantiation of OLS-OFU incorporates a self-supervised learning method associated with a loss function denoted as $l_{\rm ssl}$ for feature representation learning, and an online label shift adaptation (OLS) algorithm to effectively address distribution shift. By carefully checking the existing OLS methods and SSL methods, we identify three principles for algorithm design: maintain the theoretical guarantee, obey the underlying assumption of the existing OLS methods and fit the required condition of SSL techniques while avoiding heavy additional computational costs. Within the principles, OLS-OFU is designed as three main steps: at each time step, OLS-OFU first executes a revised OLS algorithm and then every $\tau$ steps, OLS-OFU updates the feature extractor through self-supervised learning and subsequently refines the final linear layer.
@@
In addition to its ease of implementation and seamless integration with existing OLS algorithms, OLS-OFU also shows strong theoretical guarantee and empirical performance. Theoretically, we demonstrate that OLS-OFU effectively reduces the loss of the overall algorithm by leveraging self-supervised learning (SSL) techniques to enhance the feature extractor, thereby improving predictions for test samples at each time step $t$. Empirical evaluations on various datasets, considering both online label shift and online generalized label shift scenarios, validate the effectiveness of OLS-OFU. Our OLS-OFU method achieves substantial improvements over existing OLS methods, which is as significant as the gains existing OLS methods have over the baseline (i.e., without distribution shift adaptations). This demonstrates that integrating online feature updates is as effective in solving online distribution shift as the fundamental online label shift method itself. Moreover, the improvement is consistent on various datasets, all existing OLS methods and the various choices of SSL techniques. This consistency underscoring its robustness across different scenarios and its generality to incorporate future OLS methods with more advanced online learning techniques and better SSL techniques.
@@ 2 Problem Setting & Related Work
We start with some basic notations. Let $\Delta^{K-1}$ be the probability simplex. Let $f:\mathcal{X}\to\Delta^{K-1}$ denote a classifier. Given an input $x$ from domain $\mathcal{X}$, $f(x)$ outputs a probabilistic prediction over $K$ classes. For example, $f$ can be the output from the softmax operation after any neural network. If we reweight a model output from $f$ by a vector $p\in\mathbb{R}^{K}$, we refer to this model as $g(\cdot;f,p)$ with $g$ denotes the method of reweighting. For any two vectors $p$ and $q$, $p/q$ denotes the element-wise division.
Online distribution shift adaptation. The effectiveness of any machine learning model $f$ relies on a common assumption that the train data $D_{0}$ and test data $D_{\text{test}}$ are sampled from the same distribution, i.e., $\mathcal{P}^{\rm train}=\mathcal{P}^{\rm test}$. However, this assumption is often violated in practice, which leads to distribution shift. This can be caused by various factors, such as data collection bias and changes in the data generation process. Moreover, once a well-trained model $f_{0}$ is deployed in the real world, it moves into the testing phase, which is composed of a sequence of periods or time steps. The test distribution at time step $t$, $\mathcal{P}_{t}^{\rm test}$, from which test data $x_{t}$ is sampled, may vary over time. One example is that an MRI image classifier might be trained on MRI images collected during skiing season (which may have a high frequency of head concussions) but tested afterward (when the frequency of concussion is lower). The test stage can last several months until the next classifier is trained. During this test period, the distribution of MRI images may undergo continuous changes, transitioning between non-skiing and skiing seasons.
@@
In this paper, we consider the challenging scenario where at each time step $t$, only a small batch of unlabeled samples $S_{t}=\{x_{t}^{1},\cdots,x_{t}^{B}\}$ is received. We formalize the algorithm $\mathcal{A}$ as: $\forall t\in[T]$,
@@
In contrast to the classical online learning setup, this scenario presents a significant challenge as classical online learning literature usually relies on having access to either full or partial knowledge of the loss at each time step, i.e., $\ell_{\rm sup}(f(x_{t}),y_{t})$. In this setting, however, only a batch of unlabeled samples is provided at each time step and this lack of access to label information and loss values presents a significant challenge in accurately estimating the true loss defined in Equation 1.
Online label shift (OLS) adaptation. Online label shift assumes the marginal distribution of the label $\mathcal{P}^{\rm test}_{t}(y)$ changes over time, while the conditional distribution keeps invariant:
$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y).$
@@
This assumption is most typical when the label $y$ is the causal variable and the feature $x$ is the observation  [43]. The aforementioned example of concussion detection from MRI images fits this scenario, where the presence or absence of a concussion (label) causes the observed MRI image features. Most previous methods tackle this problem through a non-trivial reduction to the classical online learning problem. Consequently, most of the online label shift algorithms [49, 7, 6] study the theoretical guarantee of the algorithm via the convergence of the regret function, which could be either static regret or dynamic regret. In previous studies, the hypothesis class $\mathcal{F}$ of the prediction function $f$ is typically chosen in one of two ways:
  1. 1.
    $\mathcal{F}$ is defined as a family of post-hoc reweightings of $f_{0}$, with the parameter space comprising reweight vectors. Examples within this category include ROGD [49], FTH [49], and FLHFTL [6].
  2. 2.
    $\mathcal{F}$ is defined as a family of functions that share the same parameters in $f_{0}$ except the last linear layer, such as UOGD [7] and ATLAS [7].
Notice that the existing OLS methods don’t update the feature extractor at the online test stage, but focus on how to leverage advanced online learning techniques to update the remaining part of the model under certain theoretical guarantees. Our method will be orthogonal to them: it will focus on how to leverage the self-supervised learning techniques to update the feature extractor and, therefore, can improve each of them. Our empirical results in fact show that the improvement from the feature extractor update is as significant as the improvement from online learning techniques.
@@
Online generalized label shift adaptation. In the context of MRI image classification, where head MRI images serve as the feature $x$, variations in software or hardware across different clinics’ MRI machines can introduce discrepancies in image characteristics like brightness, contrast, and resolution. In such scenarios, the conditional probability distribution $\mathcal{P}(x|y)$ is no longer invariant. However, when a feature extractor $h$ is robust enough, it can map images into a feature space where the images from different machines have the same distributions $\mathcal{P}(h(x)|y)$ in the transformed feature space. The concept of generalized label shift, as introduced in Tachet des Combes et al. [47], formalizes this by postulating the existence of an unknown function $h$, such that $\mathcal{P}(h(x)|y)$ remains invariant. The primary challenge in this context is to find the underlying transformation $h$. Building upon this, online generalized label shift assumes that there exists an unknown function $h$ such that
@@
Thus, to apply the OLS methods for this more general problem, it is important to learn a proper feature extractor as this underlying function $h$.
Motivations of deploying self-supervised feature updates. As we reviewed above, OLS methods in the literature don’t update the pretrained feature extractor in the online stage. However, the pretrained feature extractor in $f_{0}$ can be suboptimal for the online test stage in online (generalized) label shift adaptation, because of the three potential reasons:
  1. 1.
    The amount of pretrained data doesn’t achieve the learning capacity of the feature extractor structure, i.e. learning with more data can improve the feature extractor.
  2. 2.
    The optimal feature extractors can be different for two different distributions. In our problem, the distribution shifts overtime and the optimal feature extractor can be dynamic too.
  3. 3.
    Particularly in online generalized label shift adaptation, the domain of the data can be dramatically changed between train and test, and the pretrained feature extractor could have unpredictable performance for the test data.
Fortunately, in online (generalized) label shift adaptation problem, the learner receives many unlabeled test samples $S_{1}\cup\cdots\cup S_{t-1}$ before the prediction for time step $t$. On the other hand, self-supervised learning is very powerful for representation learning [18, 32, 30, 16, 37, 9, 23, 20, 24] and domain adaptation [46, 48, 36, 39]. In this paper, we will introduce how to deploy these self-supervised learning techniques to the existing OLS methods such that we can still enjoy the theoretical guarantees from online learning techniques and, more importantly, take the advantage of self-supervised learning to achieve better empirical performance.